# Liouville integrable Lotka-Volterra systems

### 02 Apr 2026 v1

@@ 1 Introduction
Lotka-Volterra (LV) systems are ordinary differential equations of the form
@@
where the vector $\mathbf{r}$ and matrix $\mathbf{A}$ do not depend on $\mathbf{x},t$, and $\dot{\mathbf{x}}$ denotes the $t$-derivative of $\mathbf{x}$. They are a generalisation of the 2-dimensional system, derived independently by Lotka and by Volterra [3, 14]. LV systems have been studied extensively and they found applications in e.g. biology and chemistry [1, 2, 4, 5, 6, 8, 9]. One of the interesting things about LV systems is that they can exhibit regular or chaotic behaviour, or both. Of course, systems with maximum regularity are integrable.
@@ 2 Homogeneous Liouville integrable LV systems
Integrals for (4) are functions $F$ for which $\dot{F}=\{F,H\}=0$. A Darboux Polynomial (DP) is a polynomial $P$ for which $\dot{P}=\{P,H\}=CP$, in which case $C$ is called the cofactor of $P$, also denoted $C[P]$. Clearly, the Hamiltonian itself is an integral, i.e., a DP with cofactor 0. More generally, a function $P$ which satisfies $\dot{P}=\{P,H\}=CP$ is called a second integral [1, Definition 2.14]. A product of powers of DPs, $\Pi=\prod_{i}P_{i}^{\alpha_{i}}$ is a second integral with cofactor $C[\Pi]=\sum_{i}\alpha_{i}C[P_{i}]$.
We will require linear combinations of the form $P_{I}=\alpha_{1}x_{i_{1}}+\alpha_{2}x_{i_{2}}+\cdots+\alpha_{k}x_{i_{k}}$, where $I=\{i_{1},\ldots,i_{k}\}\subset\mathbb{N}$, to be DPs. This imposes conditions on the matrix $\mathbf{A}$, as given in [7] for $k=2$, or in [10] for $k>2$. Due to the antisymmetry of $\mathbf{A}$, the coefficients $\alpha_{i}$ in $P_{I}$ are all equal; we take $\alpha_{i}=1$ for all $i$. If the matrix $\mathbf{A}$ has full rank, then each DP gives rise to a integral. Indeed, if $\mathbf{B}$ is the $1\times n$ vector such that $C[P_{I}]=\sum_{i=1}^{n}B_{i}x_{i}$, and $\mathbf{Z}=-\mathbf{B}.\mathbf{A}^{-1}$, then
@@
is an integral, cf. [7, Section 2]. Further conditions on the matrix $\mathbf{A}$ are imposed by requiring the integrals to Poisson commute. A homogeneous $n$-dimensional Hamiltonian LV system with $k$ integrals of the form (6) which are pairwise commuting will be represented by a hypergraph of order $n$ and size $k$ (i.e. a subset, with $k$ elements, of the powerset of $n$ symbols). For example, the Hamiltonian LV system with matrix
@@ 2.1.1 The even-dimensional homogeneous type $[j,0,0]$ LV system
The integrals can either by obtained from the Darboux polynomials $x_{2i-1}+x_{2i}$, $i=1,\ldots,j$, using Eq. (6), or by contraction to (n1n-1)-component systems and lifting their Casimirs, see [12, Theorem 6 and Appendix B]. We first show their cofactors vanish. Using the notation
@@ 2.1.2 The even-dimensional homogeneous type $[0,k,0]$ LV system
and $N^{n}_{l,k}=-N^{n}_{k,l}$. We denote $P_{a,b}=x_{a}+x_{a+1}+\cdots x_{b}$, following notation introduced in [12, Section 2].
@@ 2.2 Odd-dimensional homogeneous Liouville integrable LV systems
If the matrix $\mathbf{A}$ does not have full rank, Casimirs can be found by determining the null-space of $\mathbf{A}$ (which, as $\mathbf{A}$ is antisymmetric, equals the null-space of its transpose). Indeed, if $\sum_{i}A_{i,j}q_{i}=0$ for all $j$, and $\mathcal{C}=\prod_{i}x_{i}^{q_{i}}$, then
$\dot{\mathcal{C}}=\mathcal{C}\left(\sum_{i,j}q_{i}A_{i,j}x_{j}\right)=0.$
@@ 3 Nonhomogeneous Liouville integrable LV systems
A function $P(\mathbf{x})$ is homogeneous of weigth $d$ if $P(\lambda\mathbf{x})=\lambda^{d}P(\mathbf{x})$. The following theorem is a slightly improved version of [11, Theorem 3.1].
@@
Let $P(\mathbf{x})$ be a homogeneous second integral of weight $d$ with cofactor $C(\mathbf{x})$ for the system of ODEs $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$. Then $P$ is a second integral for the system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+r(\mathbf{x},t)\mathbf{x}$, with cofactor $C(\mathbf{x})+dr(\mathbf{x},t)$, where $r$ is a scalar function of $\mathbf{x},t$.
It follows that if a homogeneous LV system admits an integral $F$ which is homogeneous of weight 0, then the inhomogeneous LV system obtained by adding $\mathbf{r}=r\mathbf{x}$ to the right-hand-side admits the same integral.
@@ 3.2.1 The odd-dimensional inhomogeneous $[j,0,0]$ LV system
$R_{i}$ $=C\left(\frac{F_{i}}{F_{i+1}}\right)^{1/(a_{j+1}(c_{i}-c_{i+1}))}$
@@
We note that there is one additional DP, $P_{1,n}$, which yields an extra integral,
@@ 4.3 The even $[j,k,l]$ LV systems with $j+k+l=2$ ($n=6$)
    and commuting integrals $\left(x_{1}+x_{2}\right)^{a_{3}}x_{5}^{-c_{1}}x_{6}^{c_{1}-a_{3}}$ and $\left(x_{3}+x_{4}\right)^{a_{3}}x_{5}^{-c_{2}}x_{6}^{c_{2}-a_{3}}$. The inhomogeneous system admits an extra DP and is superintegrable.

### 27 Apr 2026 v2

@@ 1 Introduction
A large part of Jarmo Hietarinta’s academic career he has spent searching for integrable systems [4, 5, 6, 7, 8, 9, 10, 11], and one of his early interests has been Liouville integrability [3]. In this paper, we search for Liouville integrable Lotka-Volterra (LV) systems, i.e. ordinary differential equations of the form
@@
where the vector $\mathbf{r}$ and matrix $\mathbf{A}$ do not depend on $\mathbf{x},t$, and $\dot{\mathbf{x}}$ denotes the $t$-derivative of $\mathbf{x}$. They are a generalisation of the 2-dimensional system, derived independently by Lotka and by Volterra [13, 24]. So, in essence, we will be extending integrable Hamiltonian LV systems from 2 to $n$ dimensions, cf. [2].
LV systems have been studied extensively and they found applications in e.g. biology and chemistry [1, 12, 14, 15, 16, 18, 19]. One of the interesting things about LV systems is that they can exhibit regular or chaotic behaviour, or both. Of course, systems with maximum regularity are integrable.
@@ 2 Homogeneous Liouville integrable LV systems
Integrals for (4) are functions $F$ for which $\dot{F}=\{F,H\}=0$. A Darboux Polynomial (DP) is a polynomial $P$ for which
$\dot{P}=\{P,H\}=CP,$ (6)
in which case $C$ is called the cofactor of $P$, also denoted $C[P]$. Clearly, the Hamiltonian itself is an integral, i.e., a DP with cofactor 0. More generally, a function $P$ which satisfies $\dot{P}=\{P,H\}=CP$ is called a second integral [1, Definition 2.14]. A product of powers of DPs,
$\Pi=\prod_{i}P_{i}^{\alpha_{i}}$ (7)
is a second integral with cofactor $C[\Pi]=\sum_{i}\alpha_{i}C[P_{i}]$.
We will require linear combinations of the form
$P_{I}=\alpha_{1}x_{i_{1}}+\alpha_{2}x_{i_{2}}+\cdots+\alpha_{k}x_{i_{k}},$ (8)
where $I=\{i_{1},\ldots,i_{k}\}\subset\mathbb{N}_{n}=\{x_{1},\ldots,x_{n}\}$, to be DPs. This imposes conditions on the matrix $\mathbf{A}$, as given in [17] for $k=2$, or in [20] for $k>2$. Due to the antisymmetry of $\mathbf{A}$, the coefficients $\alpha_{i}$ in $P_{I}$ are all equal; we take $\alpha_{i}=1$ for all $i$. If the matrix $\mathbf{A}$ has full rank, then each DP gives rise to a integral. Indeed, if $\mathbf{B}$ is the $1\times n$ vector such that $C[P_{I}]=\sum_{i=1}^{n}B_{i}x_{i}$, and
$\mathbf{Z}=-\mathbf{B}.\mathbf{A}^{-1},$
then
@@
is an integral, cf. [17, Section 2]. Further conditions on the matrix $\mathbf{A}$ are imposed by requiring the integrals to Poisson commute. A homogeneous $n$-dimensional Hamiltonian LV system with $k$ integrals of the form (9) which are pairwise commuting will be represented by a hypergraph of order $n$ and size $k$ (i.e. a subset, with $k$ elements, of the powerset of $n$ symbols).
For example, the Hamiltonian LV system with matrix
@@ 2.1.1 The even-dimensional homogeneous type $[j,0,0]$ LV system
The integrals can either by obtained from the Darboux polynomials $x_{2i-1}+x_{2i}$, $i=1,\ldots,j$, using Eq. (9), or by contraction to (n1n-1)-component systems and lifting their Casimirs, see [22, Theorem 6 and Appendix B]. We first show their cofactors vanish. Using the notation
@@ 2.1.2 The even-dimensional homogeneous type $[0,k,0]$ LV system
and $N^{n}_{l,k}=-N^{n}_{k,l}$. We denote $P_{a,b}=x_{a}+x_{a+1}+\cdots x_{b}$, following notation introduced in [22, Section 2].
@@ 2.2 Odd-dimensional homogeneous Liouville integrable LV systems
If the matrix $\mathbf{A}$ does not have full rank, Casimirs can be found by determining the null-space of $\mathbf{A}$ (which, as $\mathbf{A}$ is antisymmetric, equals the null-space of its transpose). Indeed, if $\sum_{j}A_{i,j}q_{j}=0$ for all $i$, and $\mathcal{C}=\prod_{h}x_{h}^{q_{h}}$, then
$\{x_{i},\mathcal{C}\}=x_{i}\mathcal{C}\left(\sum_{j}A_{i,j}q_{j}\right)=0.$
@@ 3 Nonhomogeneous Liouville integrable LV systems
A function $P(\mathbf{x})$ is homogeneous of weight $d$ if $P(\lambda\mathbf{x})=\lambda^{d}P(\mathbf{x})$. The following theorem is an improved version of [21, Theorem 3.1]. Let $I=\{i_{1},\ldots,i_{k}\}\subset\mathbb{N}_{n}$, and denote $\mathbf{x}_{I}=(x_{i_{1}},\ldots,x_{i_{k}})$.
@@
Let $P(\mathbf{x}_{I})$ be a homogeneous second integral of weight $d$ with cofactor $C(\mathbf{x})$ for the system of ODEs $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})$. Then $P(\mathbf{x}_{I})$ is a second integral for the system $\dot{\mathbf{x}}=\mathbf{f}(\mathbf{x})+r(\mathbf{x},t)\mathbf{x}$, with cofactor $C(\mathbf{x})+dr(\mathbf{x},t)$, where $r$ is a scalar function of $\mathbf{x},t$.
It follows that if a homogeneous LV system admits an integral $F$ which is homogeneous of weight 0, then the inhomogeneous LV system obtained by adding $\mathbf{r}=r\mathbf{x}$ to the right-hand-side admits the same integral. Note that, in Theorem 10, the only relevant equations are the ones for $\mathbf{x}_{I}$, the equations for the complementary variables $\mathbf{x}_{I^{c}}$ can be altered differently.
@@ 3.2.1 The odd-dimensional inhomogeneous $[j,0,0]$ LV system
$R_{i}$ $=\mathcal{C}\left(\frac{F_{i}}{F_{i+1}}\right)^{1/(a_{j+1}(c_{i}-c_{i+1}))}$
@@
We note that there is one additional DP, $P_{1,n-1}$, which yields an extra integral,
@@ 4.3 The even $[j,k,l]$ LV systems with $j+k+l=2$ ($n=6$)
    and commuting integrals $\left(x_{1}+x_{2}\right)^{a_{3}}x_{5}^{-c_{1}}x_{6}^{c_{1}-a_{3}}$ and $\left(x_{3}+x_{4}\right)^{a_{3}}x_{5}^{-c_{2}}x_{6}^{c_{2}-a_{3}}$.